A 650,000 mL volume of monatomic ideal gas inside of a heat engine starts at room temperature (20.0o C) and pressure (101.325 kPa).  At first, the volume is compressed without changing the temperature.  Then, 3.7 kJ of heat is added while the volume remains constant.  Finally, the volume is allowed to expand adiabatically back to its starting conditions. Assume all processes are reversible and ideal. Prove that the entropy and efficiency for the three-stroke heat engine are consistent with the second law of thermodynamics.
Answer #1

Cycle of the three stroke heat engine is shown below +3.7 Isochore Q=0 Adiabatic - Isotherm - -- 101,325 kla VB=VC VA AB proc.. (101.325) X (10) (650000) X (10-) (10) - 27037 (8.314) X (20+273) BC process is isochoric V= constant dV = 0 W= [PdV WBC0= 0 + 3700 + AUCA AUCA = - 3700J Now Wca = - AUCA WA = 3700J In process CA TV =T.V.-1 TV, =TV- (20 + 273) X (650000) – 1 =heat absorbed by the system Here in process AB → heat rejected (Q-) BC — heat absorbed (Q+) CA → No heat exchange n= W n 94,1Entropy change in CA ASCA = A = 0 (AS)net = 0 According to second law of thermodynamics 100 percent conversion of heat into w